### v2

Abstract
We study the Bipartite Degree Realization (BDR) problem: given a graphic degree sequence $D$, decide whether it admits a realization as a bipartite graph. While bipartite realizability for a fixed vertex partition can be decided in polynomial time via the Gale–Ryser theorem, the computational complexity of BDR without a prescribed partition remains unresolved. We address this question through a parameterized analysis.
For constants $0\leq c_{1}\leq c_{2}\leq 1$, we define $\mathrm{BDR}_{c_{1},c_{2}}$ as the restriction of BDR to degree sequences of length $n$ whose degrees lie in the interval $c_{1}n,c_{2}n]$. Our main result shows that $\mathrm{BDR}_{c_{1},c_{2}}$ is solvable in polynomial time whenever $0\leq c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$, as well as for all $c_{1}>\tfrac{1}{2}$. The proof relies on a reduction to extremal least balanced degree sequences and a detailed verification of the critical Gale–Ryser inequalities, combined with a bounded subset-sum formulation.
We further show that, assuming the NP-completeness of unrestricted BDR, the problem $\mathrm{BDR}_{c_{1},c_{2}}$ remains NP-complete for all $0<c_{2}<\frac{1}{2}$ and $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$. Our results clarify the algorithmic landscape of bipartite degree realization and contribute to the broader study of potentially bipartite graphic degree sequences.
keywords:
graph realizations, bipartite graphs, degree sequences, bigraphic sequences, parameterized complexity
{MSCcodes}
05C07, 05C85, 68Q25
1 Introduction
A nonnegative integer sequence $D=(d_{1},\dots,d_{n})$ is called graphic if there exists a simple graph on $n$ vertices whose degrees are exactly $d_{i}$. We call such a graph a realization of $D$. Similarly, a pair of degree sequences $(D_{1},D_{2})$ is bigraphic if it has a bipartite graph realization, $G=(U,V,E)$, such that the degrees in $U$ are $D_{1}$ and the degrees in $V$ are $D_{2}$. In 1981, Rao considered the potentially and forcibly $\mathcal{P}$-graphic degree sequences [rao]. Given a graph property $\mathcal{P}$ and a degree sequence $D$, we say that $D$ is potentially $\mathcal{P}$-graphic if $D$ has a simple graph realization with property $\mathcal{P}$. Similarly, $D$ is forcibly $\mathcal{P}$-graphic if any realization of $D$ has property $\mathcal{P}$. For example, if the sum of the degrees in a degree sequence $D$ with length $n$ is less than 2n22n-2, then $D$ is forcibly disconnected. If the sum is at least 2n22n-2, and there is no 0 degree, then $D$ is potentially connected. If the minimum degree is greater than $\left\lfloor\frac{n-2}{2}\right\rfloor$ then $D$ is forcibly connected.
One of the open problems that Rao asked was the potentially bipartite degree sequences. According to his terminology, we call a degree sequence potentially bipartite if it admits at least one bipartite realization, and forcibly not bipartite if none of its realizations is bipartite. We will denote the potentially bipartite degree sequence problem as the BDR problem, that is, the Bipartite Degree Realization problem. The input of the BDR problem is a degree sequence $D$, and the problem is to decide if $D$ has a bipartite graph realization. This problem has an unknown computational complexity till today, although the following results are well-known. The Gale-Ryser theorem provides a polynomial time algorithm to decide if a pair of degree sequences $(D_{1},D_{2})$ has a bipartite graph realization $G=(U,V,E)$, such that the degrees in $U$ are $D_{1}$ and the degrees in $V$ are $D_{2}$ [gale, ryser]. Also, there is a polynomial running time dynamic programming algorithm to decide if a degree sequence $D$ can be split into two degree sequences, $D_{1}$ and $D_{2}$ with equal sum, a necessary condition to have a bipartite graph realization. However, there might be an exponential number of solutions how to split $D$ into two degree sequences with equal sums [barnoy1], and it is not trivial how to decide if any of them satisfies the sufficient and necessary conditions given by Gale and Ryser.
In this paper, we consider the parametric version of BDR. The input of $\textsc{BDR}_{c_{1},c_{2}}$ are degree sequences in which the degrees are between $c_{1}n$ and $c_{2}n$, where $n$ is the length of the degree sequence. The problem is still to decide if a bipartite realization exists. We have the following results. We show that $\textsc{BDR}_{c_{1},c_{2}}$ is in P if $c_{1}\leq\frac{1}{2}$ and $c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$ or $c_{1}>\frac{1}{2}$. Particularly, we show that a degree sequence $D$ of length $n$ with degrees between $c_{1}n$ and $\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}n$ is potentially bipartite if and only if $D$ can be split into two degree sequences with equal sums. Further, any degree sequence $D$ of length $n$ in which every degree is strictly greater than $\frac{n}{2}$ is necessarily forcibly not bipartite.
We also show the following. If BDR is NP-complete, then any parametric version $\textsc{BDR}_{c_{1},c_{2}}$ with $c_{1}\leq\frac{1}{2}$ and $c_{2}>\sqrt{2c_{1}}-c_{1}$ is NP-complete as well (we will show that $c_{2}>\sqrt{2c_{1}}-c_{1}$ iff $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$). This result follows from a simple polynomial-time padding (or scaling) reduction that preserves bipartite realizability.
The computational complexity of deciding whether a degree sequence admits a bipartite realization has recently been studied also by Bar-Noy et al. in a general framework that includes both simple graphs and multigraphs [barnoy1]. Our work is concerned exclusively with simple graphs. Accordingly, we restrict attention to those results of Bar-Noy et al. that apply to simple bipartite graphs and are directly comparable to ours. In this restricted setting, Bar-Noy et al. identify polynomial-time solvable cases primarily under strong structural assumptions on the degree sequence, most notably when the sequence has a high-low partition. A high-low partition of a degree sequence $D$ is a partition into $D_{1}$ and $D_{2}$ such that for all $d_{i}\in D_{1}$ and $d_{j}\in D_{2}$, $d_{i}\leq d_{j}$. A high-low partiton $(D_{1},D_{2})$ is well-behaved if the maximum degree in each vertex class is at most the size of the other vertex class. The main result of Bar-Noy et al. [barnoy1] is the following. If $D$ has a well-behaved high-low partition, then $D$ has a bipartite graph realization if and only if the well-behaved high-low partition has a bipartite graph realization. That is, if the well-behaved high-low partition does not have a bipartite graph realization then none of the equal sum partitions of the underlying degree sequence is bigraphic. These results assume a special global property (that is, well-behaved high-low partition) rather than extremal degree constraints.
In contrast, in the present work, we identify degree sequence classes in which any partition into equal sums has a bipartite graph realization. This establishes polynomial-time decidability for bipartite realizability under explicit extremal degree conditions of the form $c_{1}n\leq d_{i}\leq c_{2}n$, for constants $0\leq c_{1}\leq c_{2}<1$, covering a broad class of dense degree sequences with linear minimum and maximum degree bounds. Our results apply when a degree sequence does not have a high-low partition, and also, even when the degree sequence exhibits large variance, provided all degrees remain within the prescribed interval. The largest difference between the minimum and maximum degrees could be obtained by maximizing $\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}n-c_{1}n$ which turns out to be $(3-2\sqrt{2})n\approx 0.17157n$ when $c_{1}=\frac{3\sqrt{2}}{2}-2$. This places our tractability region in a fundamentally different part of the parameter space than that considered by Bar-Noy et al. for simple graphs.
In another work, Bar-Noy et al. also proved that the corresponding bipartite realization problem for multigraphs is NP-complete [barnoy2]. In that setting, degrees are allowed to be arbitrarily large and are not bounded as a function of the number of vertices. From this perspective, the hardness result aligns with the inherent difficulty of partitioning sequences of large integers into two parts with equal sums, a problem that is already NP-complete [garey-johnson]. Consequently, these multigraph results are not directly comparable to the simple graph setting considered here.
Bar-Noy et al. leave open the computational complexity of bipartite realizability for simple graphs when degrees are bounded only by n1n-1 [barnoy2]. Our results contribute to this open direction by showing that if bipartite realizability for simple graphs with degrees bounded by n1n-1 were NP-complete, then NP-completeness would persist even under much stronger restrictions, namely for any upper bound of the form $cn$ for any fixed constant $c>0$.
2 Preliminaries
We introduce the necessary ingredients to achieve the stated results. We are going to change graphs using hinge-flips. A hinge-flip operation removes an edge $(u_{1},v)$ and add an edge $(u_{2},v)$ not presented in the edge set before the operation. The hinge-flip operations do not change the sum of the degrees, however, they do change the degree sequence of the graph on which they act. However, observe the following: if in a pair of degree sequences $(D_{1},D_{2})$ the degree of $u_{i}$ is greater than the degree of $u_{j}$, then in any realization of $(D_{1},D_{2})$ there exists a hinge-flip that removes an edge $(u_{i},v)$ and adds an edge $(u_{j},v)$. That is, if a pair of degree sequences $(D_{1},D_{2})$ is bigraphic than any pair of degree sequences $(D_{1}^{\prime},D_{2})$ is also bigraphic that can be obtained from $(D_{1},D_{2})$ by decreasing a degree $d_{j}\in D_{1}$ and increasing a $d_{i}\in D_{1}$ with $d_{i}<d_{j}$. When $d_{i}<d_{j}-1$, we call the corresponding hinge-flip a balancing hinge-flip (if $d_{i}=d_{j}+1$, the hinge flip only swap the two degrees). Observe that a balancing hinge-flip does not change the sum of the degrees, cannot decrease the minimum degree as well as cannot increase the maximum degree. This observation leads to the notation of the least balanced degree sequence and a lemma on it.
Definition 2.1.
Let $\mathcal{D}(n,\Sigma,d,\Delta)$ denote the set of degree sequences of length $n$ with the sum of degrees is equal to $\Sigma$ such that the smallest degree is greater than or equal to $d$ and the largest degree is smaller than or equal to $\Delta$. The least balanced degree sequence in this set is
$LDBS(n,\Sigma,\Delta,d):=(\underbrace{\Delta,\Delta,\ldots,\Delta}_{k},d_{int},\underbrace{d,d,\ldots,d}_{n-k-1})$
where $k=\left\lfloor\frac{\Sigma-nd}{\Delta-d}\right\rfloor$ and $d_{int}=\Sigma-k\Delta-(n-k-1)d$.
We have the following lemma.
Lemma 2.2.
Let $(D_{1},D_{2})$ be a pair of degree sequences. If $D_{1}\in\mathcal{D}(n_{1},\Sigma,\Delta_{1},d_{1})$, $D_{2}\in\mathcal{D}(n_{2},\Sigma,\Delta_{2},d_{2})$ and $(LBDS(n_{1},\Sigma,\Delta_{1},d_{1}),LBDS(n_{2},\Sigma,\Delta_{2},d_{2}))$ is bigraphic then $(D_{1},D_{2})$ is also bigraphic.
The proof can be found in [limiklos]; an analogous observation on simple degree sequences was already observed by Aigner and Triesch [at94].
The graphicality of a bipartite degree sequence can be checked with the Gale-Ryser theorem [gale, ryser]. The Gale-Ryser theorem says that a bipartite degree sequence $D:=(D_{1},D_{2})$ with $D_{1}=d_{1,1}\geq d_{1,2}\geq\ldots\geq d_{1,n}$ and $D_{2}=d_{2,1},d_{2,2},\ldots,d_{2,m}$ is bigraphic if and only if the sum of the degrees in $D_{1}$ equals the sum of the degrees in $D_{2}$ and for all $k=1,2,\ldots,n$,
(1) $\sum_{i=1}^{k}d_{1,i}\leq\sum_{j=1}^{m}\min\{k,d_{2,j}\}.$
We make the following observation. A bipartite degree sequence $D:=(D_{1},D_{2})$ with $D_{1}=d_{1,1}\geq d_{1,2}\geq\ldots\geq d_{1,n}$ and $D_{2}=d_{2,1},d_{2,2},\ldots,d_{2,m}$ is bigraphic if and only if the degree sequence $D^{\prime}:=d_{1,1}+n-1,d_{1,2}+n-1,\ldots,d_{1,n}+n-1,d_{2,1},d_{2,2},\ldots,d_{2,m}:=$ $d^{\prime}_{1},d^{\prime}_{2},\ldots d^{\prime}_{n^{\prime}}$ is graphic. Then for all $k\leq n$, the Gale-Ryser inequalities in equation 1 are equivalent with the Erdős-Gallai inequalities [eg]:
(2) $\sum_{i=1}^{k}d^{\prime}_{i}\leq k(k-1)+\sum_{j=k+1}^{n^{\prime}}\min\{k,d_{j}\}.$
Further, the Erdős-Gallai inequalities naturally hold for all $k>n$ for $D^{\prime}$ if for all $j$, $d_{2,j}\leq n$.
Tripathi and Vijay [TV03] proved that a degree sequence is graphic if and only if the sum of the degrees is even and for all $k$ with $d_{k}>d_{k+1}$ the Erdős-Gallai inequalities hold. It leads to the following observation. {observation}
The bipartite degree sequence
$D=(LDBS(n_{1},\Sigma,\Delta_{1},d_{1}),LDBS(n_{2},\Sigma,\Delta_{2},d_{2}))$
is bigraphic if and only if the Gale-Ryser inequality with $k=\left\lfloor\frac{\Sigma-n_{1}d_{1}}{\Delta_{1}-d_{1}}\right\rfloor$ and $k=\left\lceil\frac{\Sigma-n_{1}d_{1}}{\Delta_{1}-d_{1}}\right\rceil$ are satisfied.
3 Polynomial algorithm for $\textsc{BDR}_{c_{1},c_{2}}$
We are going to prove the tractable result of the parameterized BDR problem. Before we state and prove the theorem on the tractable result, we need a technical observation. {observation} Let $c_{1}\geq 0$ and $0\leq c_{2}<1$. Then
$c_{1}\;\geq\;\frac{c_{2}^{2}}{1-c_{2}}\quad\text{if and only if}\quad c_{2}\;\leq\;\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}.$
Proof 3.1.
Since $0\leq c_{2}<1$, the denominator $1-c_{2}$ is positive. Therefore,
$c_{1}\geq\frac{c_{2}^{2}}{1-c_{2}}$
is equivalent to
(3) $c_{1}(1-c_{2})\geq c_{2}^{2}.$
Rearranging (3), we obtain
$c_{2}^{2}+c_{1}c_{2}-c_{1}\leq 0.$
We now view this as a quadratic inequality in the variable $c_{2}$. The corresponding quadratic equation
$c_{2}^{2}+c_{1}c_{2}-c_{1}=0$
has discriminant
$\Delta=c_{1}^{2}+4c_{1}=c_{1}(c_{1}+4).$
Thus, the two real roots are
$c_{2}=\frac{-c_{1}\pm\sqrt{c_{1}(c_{1}+4)}}{2}.$
Since $c_{2}\geq 0$, only the positive root is relevant. Therefore,
$c_{2}^{2}+c_{1}c_{2}-c_{1}\leq 0\quad\Longleftrightarrow\quad 0\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}.$
This establishes the claimed equivalence.
Theorem 3.2.
Let $c_{1}$ and $c_{2}$ be constants satisfying $0\leq c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$. Then the $\textsc{BDR}_{c_{1},c_{2}}$ is in P. Also, for all $\frac{1}{2}<c_{1}\leq c_{2}<1$, the $\textsc{BDR}_{c_{1},c_{2}}$ is in P.
Proof 3.3.
The case when $c_{1}>\frac{1}{2}$ is trivial: if the minimum degree is larger than $\frac{n}{2}$ in a degree sequence $d$ of length $n$, then $D$ cannot have a bipartite graph realization. Indeed, in any bipartite graph $G=(U,V,E)$ with $|U|+|V|=n$, at least one of $|U|$ and $|V|$ is at most $\frac{n}{2}$. Then any degree in the opposite vertex class must have degree at most $\frac{n}{2}$.
Observe the following. The inequality $c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$ implies that $c_{1}\leq\frac{1}{2}$ because for any $c_{1}>\frac{1}{2}$ it holds that $c_{1}>\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$. When $0\leq c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$ , it is enough to show that for any $s_{1}\leq s_{2}$ with $s_{1}+s_{2}=n$, and for any $\Sigma$ with $c_{1}s_{2}\leq\Sigma\leq c_{2}s_{1}$, the bipartite degree sequence $D:=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}n,c_{2}n))$ is graphic. Indeed, according to Lemma 2.2, then any degree sequence $D^{\prime}:=(D_{1},D_{2})$ with $D_{1}\in\mathcal{D}(s_{1},\Sigma,c_{1}n,c_{2}n)$ and $D_{2}\in\mathcal{D}(s_{2},\Sigma,c_{1}n,c_{2}n)$ is graphic. Therefore, a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ has a bipartite graph realization if and only if $D$ has a split of two degree sequences with equal sum. This is the subset sum problem, which is NP-complete in general, but has a pseudo-polynomial time algorithm. The pseudopolinomiality means that the runtime of the algorithm grows polynomial with the values of the numbers and not the number of digits needed to describe the numbers. However, any degree is at most $\frac{n}{2}$, therefore the subset sum problem is actually in P for these degree sequences.
We use the Gale-Ryser theorem to prove that the bipartite degree sequence $D$ is graphic. According to Observation 2, it is enough to check the Gale-Ryser inequalities with $k=\left\lfloor\frac{\Sigma-s_{1}c_{1}n}{(c_{2}-c_{1})n}\right\rfloor$ and $k=\left\lceil\frac{\Sigma-s_{1}c_{1}n}{(c_{2}-c_{1})}\right\rceil$.
With these considerations, we assume the nonnegative parameters satisfy
$0<c_{1}<c_{2}<\frac{1}{2},$
together with the condition
$c_{1}\geq\frac{c_{2}^{2}}{1-c_{2}},$
which is equivalent with $c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$, according to Observation 3. There are possible intermediate degrees with
$c_{1}n\leq d_{1}\leq c_{2}n,\qquad c_{1}n\leq d_{2}\leq c_{2}n,$
and we assume that the sums of the degrees in the two degree sequences are the same, that is,
(4) $m_{1}c_{1}n+d_{1}+n_{1}c_{2}n\;=\;m_{2}c_{1}n+d_{2}+n_{2}c_{2}n,$
We also know that the number of different degrees satisfy the equation
$m_{1}+m_{2}+n_{1}+n_{2}+2=n.$
By symmetry between the two vertex classes, we may assume without loss of generality that $n_{1}+m_{1}\geq n_{2}+m_{2}$.
We have to check those Gale-Ryser inequalities which were obtained in Observation 2. The first is:
(5) $n_{1}c_{2}n\;\leq\;n_{2}\min\{n_{1},c_{2}n\}\;+\;\min\{n_{1},d_{2}\}\;+\;m_{2}\min\{n_{1},c_{1}n\}.$
We now prove equation 5 by splitting into three exhaustive cases according to the relative size of $n_{1}$.
Case (A). $n_{1}\geq c_{2}n$.
In this range we have $\min\{n_{1},c_{2}n\}=c_{2}n$, $\min\{n_{1},d_{2}\}=d_{2}$ (because $d_{2}<c_{2}n<n_{1}$) and $\min\{n_{1},c_{1}n\}=c_{1}n$. Thus, the right-hand side of equation 5 equals
$n_{2}c_{2}n+d_{2}+m_{2}c_{1}n.$
Using equation 4,
$n_{1}c_{2}n\leq n_{1}c_{2}n+d_{1}+m_{1}c_{1}n=n_{2}c_{2}n+d_{2}+m_{2}c_{1}n.$
This is exactly the desired inequality in Case (A). Hence equation 5 holds when $n_{1}\geq c_{2}n$.
Case (B). $n_{1}\leq c_{1}n$.
Here $\min\{n_{1},c_{2}n\}=n_{1}$, $\min\{n_{1},d_{2}\}=n_{1}$ (since $d_{2}\geq c_{1}n\geq n_{1}$), and $\min\{n_{1},c_{1}n\}=n_{1}$. Therefore the right-hand side of equation 5 equals
$n_{2}n_{1}+n_{1}+m_{2}n_{1}=(n_{2}+m_{2}+1)n_{1}.$
So equation 5 becomes
$n_{1}c_{2}n\leq(n_{2}+m_{2}+1)n_{1}.$
If $n_{1}=0$ this inequality is trivial. Otherwise, cancel $n_{1}>0$ to obtain the equivalent condition
(6) $c_{2}\leq\frac{n_{2}+m_{2}+1}{n}.$
We claim that the smallest possible value for $\frac{n_{2}+m_{1}+1}{n}$ is $\frac{c_{1}}{c_{2}+c_{1}}$. Indeed, the smallest fraction is obtained when the longer degree sequence contains only $c_{1}n$ degrees, and the shorter degree sequence contains only $c_{2}n$ degrees. That is, $m_{2}=0$, $d_{2}=c_{2}n$, $n_{1}=0$ and $d_{1}=c_{1}n$. Denote $x:=n_{2}+m_{1}+1$, and then we obtain $c_{1}(n-x)=c_{2}x$, that is, $x=\frac{c_{1}n}{c_{2}+c_{1}}$. Therefore
$\frac{c_{1}}{c_{2}+c_{1}}\leq\frac{n_{1}+m_{1}+1}{n}.$
However, we know that $c_{2}\leq\frac{c_{1}}{c_{2}+c_{1}}$ for it is equivalent with the condition $\frac{c_{2}^{2}}{1-c_{2}}\leq c_{1}$. Therefore, equation 5 holds in Case (B).
Case (C). $c_{1}n<n_{1}<c_{2}n$.
In this range we have $\min\{n_{1},c_{2}n\}=n_{1}$, and $\min\{n_{1},c_{1}n\}=c_{1}n$. Hence equation 5 becomes
$n_{1}c_{2}n\leq n_{2}n_{1}+\min\{n_{1},d_{2}\}+m_{2}c_{1}n.$
We introduce auxiliary variables $x$ and $y$ measuring the deviation from a balanced split of the degree sequence containing equal number of $c_{1}n$ and $c_{2}n$ degrees. Using these new variables, we rewrite equation 4 in the following form:
(7) $n_{1}c_{2}n+d_{1}+\left(\frac{n}{2}-n_{1}+x-1\right)c_{1}n=(n_{1}+y)c_{2}n+d_{2}+\left(\frac{n}{2}-n_{1}-x-y-1\right)c_{1}n.$
From this, we obtain
$y=\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}.$
Then the Gale-Ryser inequality will be in form
$n_{1}c_{2}n\leq\left(n_{1}+\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}\right)n_{1}+\min\{n_{1},d_{2}\}+$
$\left(\frac{n}{2}-n_{1}-x-\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}-1\right)c_{1}n.$
To simplify this equation, we divide it by $n^{2}$, and introduce new variables $\eta:=\frac{n_{1}}{n}$, $z:=\frac{x}{n}$. After rearranging, we get
$0$ $\leq$ $\underbrace{\eta^{2}-(c_{1}+c_{2})\eta+\frac{c_{1}}{2}}_{R_{1}}+\underbrace{\frac{\frac{d_{1}-d_{2}}{n^{2}}}{c_{2}-c_{1}}(\eta-c_{1})+\frac{\min\{\eta,\frac{d_{2}}{n}\}}{n}}_{R_{2}}+$
$+\underbrace{z\left(\frac{2c_{1}}{c_{2}-c_{1}}\eta-\left(1+\frac{2c_{1}}{c_{2}-c_{1}}\right)c_{1}\right)}_{R_{3}}.$
We will analyze $R_{1}$, $R_{2}$ and then conclude that the inequality holds for all $\eta$ and $z$. $R_{1}$ can be written as
$\left(\eta-\frac{c_{1}+c_{2}}{2}\right)^{2}-\left(\frac{c_{1}+c_{2}}{2}\right)^{2}+\frac{c_{1}}{2}.$
Applying $c_{1}=\frac{c_{2}^{2}}{1-c_{2}}$, it follows immediately that
$\left(\frac{c_{1}+c_{2}}{2}\right)^{2}+\frac{c_{1}}{2}=-\left(\frac{\frac{c_{2}^{2}}{1-c_{2}}+c_{2}}{2}\right)^{2}+\frac{\frac{c_{2}^{2}}{1-c_{2}}}{2}=\frac{(c_{2}^{2}-2c_{2}^{3})}{(4(c_{2}-1)^{2}}$
is positive for all $0<c_{2}<\frac{1}{2}$. Therefore $R_{1}$ is always positive.
The part $R_{2}$ can be written as
$\frac{\frac{\frac{d_{1}-d_{2}}{n}}{c_{2}-c_{1}}(\eta-c_{1})+\min\{\eta,\frac{d_{2}}{n}\}}{n}.$
Recall that $\eta>c_{1}$ for $n_{1}>c_{1}n$. If $d_{1}-d_{2}\geq 0$, then we can immediately see that $R_{2}$ is positive. Now assume that $d_{1}-d_{2}<0$. Then observe that $\left|\frac{d_{1}-d_{2}}{n}\right|\leq c_{2}-c_{1}$, and also, $\eta-c_{1}\leq c_{2}-c_{1}$ for $n_{1}\leq c_{2}n$. Now if $\min\{\eta,\frac{d_{2}}{n}\}=\eta$, then we get
$\frac{\frac{\frac{d_{1}-d_{2}}{n}}{c_{2}-c_{1}}(\eta-c_{1})+\min\{\eta,\frac{d_{2}}{n}\}}{n}\geq\frac{c_{1}-\eta+\eta}{n}>0.$
On the other hand, if $\min\{\eta,\frac{d_{2}}{n}\}=\frac{d_{2}}{n}$, then
$\frac{\frac{\frac{d_{1}-d_{2}}{n}}{c_{2}-c_{1}}(\eta-c_{1})+\min\{\eta,\frac{d_{2}}{n}\}}{n}\geq\frac{\frac{d_{1}-d_{2}}{n}+\frac{d_{2}}{n}}{n}>0.$
Therefore $R_{2}$ is always positive. Therefore, we can look at equation 3.3 as
$0\leq a+zb=f(z),$
where $a$ is positive. The linear function $f(z)$ is positive on the interval $0,\max\{z\}]$, if it is positive at the endpoints of the interval.(Recall that we assume that $n_{1}+m_{1}\geq n_{2}+m_{2}$, therefore $z\geq 0$.) We know that it is positive when $z=0$ for $a$ is positive. When $z$ takes its maximum, then the shorter degree sequence contains only $c_{2}n$ values and at most one $d_{2}$. However, we already have shown that in that case the number of $c_{2}n$ degrees is at least $c_{2}n$, therefore, the Gale-Ryser inequality is satisfied. This is equivalent with $f(z)$ is non-negative at the $\max\{z\}$ endpoint. It follows that the inequality in equation 3.3 always holds.
Now we turn to the second Gale-Ryser inequality, that is,
(9) $n_{1}c_{2}n+d_{1}\;\leq\;n_{2}\min\{n_{1}+1,c_{2}n\}\;+\;\min\{n_{1}+1,d_{2}\}\;+\;m_{2}\min\{n_{1}+1,c_{1}n\},$
under the standing assumptions and with the notation of the theorem. Again, we prove equation 9 by splitting into three exhaustive cases according to the relative size of $n_{1}$.
Case (A): $n_{1}+1\geq c_{2}n$.
Here $\min\{n_{1}+1,c_{2}n\}=c_{2}n$. Also $d_{2}<c_{2}n\leq n_{1}+1$, so $\min\{n_{1}+1,d_{2}\}=d_{2}$. Finally, $n_{1}+1\geq c_{2}n>c_{1}n$, hence $\min\{n_{1}+1,c_{1}n\}=c_{1}n$. Therefore the right-hand side of equation 9 equals
$n_{2}c_{2}n+d_{2}+m_{2}c_{1}n.$
Just like in the case of the first critical Gale-Ryser inequality, applying equation 4, we get
$n_{1}c_{2}n+d_{1}\leq n_{1}c_{2}n+d_{1}+m_{1}c_{1}n=n_{2}c_{2}n+d_{2}+m_{2}c_{1}n.$
Thus equation 9 holds in Case (A).
Case (B). $n_{1}+1\leq c_{1}n$.
Now $\min\{n_{1}+1,c_{2}n\}=n_{1}+1$, $\min\{n_{1}+1,d_{2}\}=n_{1}+1$ (because $d_{2}\geq c_{1}n\geq n_{1}+1$), and $\min\{n_{1}+1,c_{1}n\}=n_{1}+1$. Hence the right-hand side equals
$n_{2}(n_{1}+1)+(n_{1}+1)+m_{2}(n_{1}+1)=(n_{2}+m_{2}+1)(n_{1}+1).$
We must show
$n_{1}c_{2}n+d_{1}\leq(n_{2}+m_{2}+1)(n_{1}+1).$
Use the trivial upper bound $d_{1}\leq c_{2}n$ on $d_{1}$ to get
$n_{1}c_{2}n+d_{1}\leq(n_{1}+1)c_{2}n.$
Therefore it suffices to prove
$(n_{1}+1)c_{2}n\leq(n_{2}+m_{2}+1)(n_{1}+1),$
which reduces to
(B’) $c_{2}n\leq n_{2}+m_{2}+1.$
As in the proof of the first Gale–Ryser inequality, the minimal possible value of $n_{2}+m_{2}+1$ (subject to the total-count constraint $m_{1}+m_{2}+n_{1}+n_{2}+2=n$ and the degree-sum bounds) is attained when the longer side is filled with the small degrees and the shorter side with the large degrees. One checks (exactly as in Case (B) of the first inequality) that this gives the lower bound
$\frac{n_{2}+m_{2}+1}{n}\geq\frac{c_{1}}{c_{1}+c_{2}},$
so (B’) follows from the parameter condition
$c_{2}\leq\frac{c_{1}}{c_{1}+c_{2}},$
which is equivalent to $c_{1}\geq\dfrac{c_{2}^{2}}{1-c_{2}}$ (the hypothesis of the theorem). Thus equation 9 holds in Case (B).
Case (C). $c_{1}n<n_{1}+1<c_{2}n$. In this case we have
$\min\{n_{1}+1,c_{2}n\}=n_{1}+1,\qquad\min\{n_{1}+1,c_{1}n\}=c_{1}n,$
We again use equation 7 to get the Gale-Ryser inequality
$n_{1}c_{2}n+d_{1}$ $\leq$ $\left(n_{1}+\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}\right)(n_{1}+1)+$
$+\min\{n_{1}+1,d_{2}\}+\left(\frac{n}{2}-x-\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}-1\right)c_{1}n.$
Again, dividing by $n^{2}$, introducing new variables $\eta=\frac{n_{1}}{n}$, $z=\frac{x}{n}$ and rearranging the inequality we get
$0$ $\leq$ $\underbrace{\eta^{2}-(c_{1}+c_{2})\eta+\frac{c_{1}}{2}+\frac{\eta}{n}}_{R_{1}}+\underbrace{\frac{\frac{d_{1}-d_{2}}{n^{2}}}{c_{2}-c_{1}}\left(\eta+\frac{1}{n}-c_{1}\right)+\frac{\min\left\{\eta+\frac{1}{n},\frac{d_{2}}{n}\right\}}{n}}_{R_{2}}+$
$+\underbrace{z\left(\frac{2c_{1}}{c_{2}-c_{1}}\left(\eta+\frac{1}{n}\right)-\left(1+\frac{2c_{1}}{c_{2}-c_{1}}\right)c_{1}\right)}_{R_{3}}.$
The expression $R_{1}$ in equation 3.3 is increased by a positive term ($\frac{\eta}{n}$) compared to the expression $R_{1}$ in equation 3.3, therefore it is positive. $\eta+\frac{1}{n}$ in expression $R_{2}$ in equation 3.3 plays the same role than $\eta$ in expression $R_{2}$ in equation 3.3, therefore $R_{2}$ is positive. That is, the right hand side is again a linear function $a+bz$ with $a>0$. Again, it is non-negative on the whole interval $0,\max\{z\}]$ if it is non-negative at the endpoints. It is positive when $z=0$ for $a>0$, and when $z$ has maximum, the shorter side of the bipartite degree sequence contains at least $c_{2}n$ number of $c_{2}n$ degrees, therefore, the Gale-Ryser inequality holds, therefore $a+bz$ is non-negative when $z$ takes its maximum. Therefore, the second critical Gale-Ryser inequality also holds, therefore, the degree sequence $D=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}nc_{2}n))$ is graphic.
We proved that for any $s_{1}\leq s_{2}$ with $s_{1}+s_{2}=n$, and for any $\Sigma$ with $c_{1}s_{2}\leq\Sigma\leq c_{2}s_{1}$, the bipartite degree sequence
$D:=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}n,c_{2}n))$
is bigraphic. Therefore, a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ has a bipartite graph realization if and only if it can be partitioned into two degree sequences with equal sum. Since it can be checked in polynomial time if a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ can be split into two degree sequences with equal sums, the $\textsc{BDS}_{c_{1},c_{2}}$ is in P for all $0<c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$.
4 The conditional NP-completeness result
Now we turn to the other main result of the paper. Before we prove the main theorem, we state and prove a bunch of technical observations. They do not go beyond high school level mathematics, but on the other hand, they are not trivial, and we give formal proof for the sake of completeness. {observation} For $0<c_{1}<c_{2}<\frac{1}{2}$,
$c_{1}<1-c_{2}-\sqrt{1-c_{2}}\,\,\Leftrightarrow\,\,c_{2}>\sqrt{2c_{1}}-c_{1}.$
Proof 4.1.
Let $0<c_{1}<c_{2}<\tfrac{1}{2}$.
We first prove the forward implication.
Assume
$c_{1}<1-c_{2}-\sqrt{1-c_{2}}.$
Since $c_{2}<\frac{1}{2}$, we have $1-c_{2}>0$, and hence the square root is well-defined. Rewriting the inequality gives
$\sqrt{1-c_{2}}<1-c_{2}-c_{1}.$
Both sides are nonnegative because
$1-c_{2}-c_{1}>1-2c_{2}>0.$
Squaring both sides preserves the inequality:
$1-c_{2}<(1-c_{2}-c_{1})^{2}.$
Expanding the right-hand side yields
$1-c_{2}<1-2(c_{1}+c_{2})+(c_{1}+c_{2})^{2}.$
Canceling $1$ from both sides and rearranging gives
$0<c_{2}^{2}+2c_{1}c_{2}+c_{1}^{2}-2c_{1}-c_{2}.$
This simplifies to
$c_{2}^{2}+(2c_{1}-1)c_{2}+(c_{1}^{2}-2c_{1})>0.$
The quadratic expression in $c_{2}$ has discriminant
$(2c_{1}-1)^{2}-4(c_{1}^{2}-2c_{1})=1+4c_{1}.$
Its roots are
$c_{2}=\sqrt{2c_{1}}-c_{1}\quad\text{and}\quad c_{2}=-\sqrt{2c_{1}}-c_{1}.$
Since $c_{2}>0$, the inequality holds if and only if
$c_{2}>\sqrt{2c_{1}}-c_{1}.$
Now we prove the reverse implication. Assume
$c_{2}>\sqrt{2c_{1}}-c_{1}.$
Reversing the algebraic steps above shows that
$c_{2}^{2}+(2c_{1}-1)c_{2}+(c_{1}^{2}-2c_{1})>0,$
which implies
$1-c_{2}<(1-c_{2}-c_{1})^{2}.$
Since both sides are nonnegative, taking square roots yields
$\sqrt{1-c_{2}}<1-c_{2}-c_{1},$
or equivalently,
$c_{1}<1-c_{2}-\sqrt{1-c_{2}}.$
Therefore,
$c_{1}<1-c_{2}-\sqrt{1-c_{2}}\quad\text{if and only if}\quad c_{2}>\sqrt{2c_{1}}-c_{1}.$
{observation}
For $0<c_{1}<c_{2}<\frac{1}{2}$,
$\frac{c_{1}+c_{2}}{2}\frac{c_{2}-c_{1}}{2}=\frac{1-c_{2}-c_{1}}{2}c_{1}\,\,\Leftrightarrow c_{1}=1-c_{2}-\sqrt{1-2c_{2}}$
Proof 4.2.
Assume $0<c_{1}<c_{2}<\tfrac{1}{2}$.
We first prove the forward implication. Suppose
$\frac{c_{1}+c_{2}}{2}\frac{c_{2}-c_{1}}{2}=\frac{1-c_{2}-c_{1}}{2}c_{1}.$
Multiplying both sides by $4$ gives
$(c_{1}+c_{2})(c_{2}-c_{1})=2c_{1}(1-c_{2}-c_{1}).$
Expanding both sides yields
$c_{2}^{2}-c_{1}^{2}=2c_{1}-2c_{1}c_{2}-2c_{1}^{2}.$
Rearranging terms, we obtain
$0=c_{1}^{2}+c_{2}^{2}+2c_{1}c_{2}-2c_{1}.$
This can be written as
$(c_{1}+c_{2})^{2}=2c_{1}.$
Viewing this as a quadratic equation in $c_{1}$, we rewrite it as
$c_{1}^{2}+2(c_{2}-1)c_{1}+c_{2}^{2}=0.$
The discriminant of this quadratic is
$4(c_{2}-1)^{2}-4c_{2}^{2}=4(1-2c_{2}).$
Since $c_{2}<\tfrac{1}{2}$, the discriminant is positive. Solving for $c_{1}$ yields
$c_{1}=\frac{2(1-c_{2})\pm 2\sqrt{1-2c_{2}}}{2}=(1-c_{2})\pm\sqrt{1-2c_{2}}.$
Because $c_{1}<c_{2}$, the plus sign is impossible. Therefore,
$c_{1}=1-c_{2}-\sqrt{1-2c_{2}}.$
Now we prove the reverse implication. Assume
$c_{1}=1-c_{2}-\sqrt{1-2c_{2}}.$
Then
$(c_{1}+c_{2})^{2}=2c_{1},$
and reversing the algebraic steps above yields
$(c_{1}+c_{2})(c_{2}-c_{1})=2c_{1}(1-c_{2}-c_{1}).$
Dividing both sides by $4$ gives
$\frac{c_{1}+c_{2}}{2}\cdot\frac{c_{2}-c_{1}}{2}=\frac{1-c_{2}-c_{1}}{2}\cdot c_{1}.$
Hence,
$\frac{c_{1}+c_{2}}{2}\cdot\frac{c_{2}-c_{1}}{2}=\frac{1-c_{2}-c_{1}}{2}\cdot c_{1}\;\;\Longleftrightarrow\;\;c_{1}=1-c_{2}-\sqrt{1-2c_{2}}.$
{observation}
$\left(\frac{c_{1}+c_{2}}{2}\right)^{2}+\frac{1-c_{1}-c_{2}}{2}c_{1}=\frac{c_{1}+c_{2}}{2}c_{2}\,\,\Leftrightarrow\,\,c_{2}=\sqrt{2c_{1}}-c_{1}$
Proof 4.3.
Assume $0<c_{1}<c_{2}<\tfrac{1}{2}$. We prove the equivalence. First, suppose
$\left(\frac{c_{1}+c_{2}}{2}\right)^{2}+\frac{1-c_{1}-c_{2}}{2}\cdot c_{1}=\frac{c_{1}+c_{2}}{2}\cdot c_{2}.$
Expanding the left-hand side gives
$\left(\frac{c_{1}+c_{2}}{2}\right)^{2}=\frac{(c_{1}+c_{2})^{2}}{4}=\frac{c_{1}^{2}+2c_{1}c_{2}+c_{2}^{2}}{4},$
and
$\frac{1-c_{1}-c_{2}}{2}\cdot c_{1}=\frac{c_{1}-c_{1}^{2}-c_{1}c_{2}}{2}=\frac{2c_{1}-2c_{1}^{2}-2c_{1}c_{2}}{4}.$
Adding these terms yields
$\frac{-c_{1}^{2}+c_{2}^{2}+2c_{1}}{4}.$
The right-hand side expands to
$\frac{c_{1}+c_{2}}{2}\cdot c_{2}=\frac{c_{1}c_{2}+c_{2}^{2}}{2}=\frac{2c_{1}c_{2}+2c_{2}^{2}}{4}.$
Hence the equality holds if and only if
$c_{1}^{2}+c_{2}^{2}+2c_{1}=2c_{1}c_{2}+2c_{2}^{2}.$
Rearranging terms gives
$0=c_{1}^{2}+2c_{1}c_{2}+c_{2}^{2}-2c_{1},$
or equivalently,
$(c_{1}+c_{2})^{2}=2c_{1}.$
We now solve this equation for $c_{2}$. Writing it as a quadratic equation in $c_{2}$, we obtain
$c_{2}^{2}+2c_{1}c_{2}+(c_{1}^{2}-2c_{1})=0.$
The discriminant is
$(2c_{1})^{2}-4(c_{1}^{2}-2c_{1})=8c_{1}.$
Thus,
$c_{2}=\frac{-2c_{1}\pm 2\sqrt{2c_{1}}}{2}=-c_{1}\pm\sqrt{2c_{1}}.$
Since $c_{2}>0$, only the positive root is admissible, and hence
$c_{2}=\sqrt{2c_{1}}-c_{1}.$
Conversely, assume
$c_{2}=\sqrt{2c_{1}}-c_{1}.$
Then
$(c_{1}+c_{2})^{2}=2c_{1},$
and reversing the algebraic steps above yields
$\left(\frac{c_{1}+c_{2}}{2}\right)^{2}+\frac{1-c_{1}-c_{2}}{2}\cdot c_{1}=\frac{c_{1}+c_{2}}{2}\cdot c_{2}.$
Therefore,
$\left(\frac{c_{1}+c_{2}}{2}\right)^{2}+\frac{1-c_{1}-c_{2}}{2}\cdot c_{1}=\frac{c_{1}+c_{2}}{2}\cdot c_{2}\;\;\Longleftrightarrow\;\;c_{2}=\sqrt{2c_{1}}-c_{1}.$
We do not know whether or not BDR is NP-complete. However, we can prove the following theorem.
Theorem 4.4.
If BDR is NP-complete, then $\textsc{BDR}_{c_{1},c_{2}}$ is also NP-complete for all $0<c_{2}<\frac{1}{2}$ and $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$.
Proof 4.5.
We use polynomial reduction in the following way. Given $c_{1}$ and $c_{2}$ parameters with the prescribed constraints, for any degree sequence $D$ coming from the BDR problem, we are going to construct a degree sequences $D^{\prime}$ in polynomial time, such that $D$ has a bipartite graph realization if and only if $D^{\prime}$ has a bipartite graph realization.
First we need a technical computation that depends only on $c_{1}$ and $c_{2}$ and not on $D$ (therefore, it can be done in constant time). Recall that $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$ is equivalent with $c_{2}>\sqrt{2c_{1}}-c_{1}$, according to Observation 4. Let $c_{2}^{\prime}:=\sqrt{2c_{1}}-c_{1}$. Since $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$, it follows that $c_{2}^{\prime}<c_{2}$. Let $r$ be a rational number between $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ obtained in the following way: in the decimal representation of $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ let $k$ be the first digit in which the said two numbers differ. Take $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ rounded to $k$ digits (these are rational numbers) and take the arithmetic average of them. Let $r$ be the so obtained rational number. Then $c_{2}^{\prime}<\frac{1-r^{2}}{2}<c_{2}$. Since the function $1-x-\sqrt{1-2x}$ is strictly increasing on the interval $0,\frac{1}{2}]$, it follows that $c_{1}<\frac{(1-r)^{2}}{2}\left(=1-\frac{1-r^{2}}{1}-\sqrt{1-2\frac{1-r^{2}}{2}}\right)<1-c_{2}-\sqrt{1-2c_{2}}$. For short, we define $\tilde{c_{1}}:=\frac{(1-r)^{2}}{2}$ and $\tilde{c_{2}}:=\frac{1-r^{2}}{2}$. Since both $\tilde{c_{1}}$ and $\tilde{c_{2}}$ are rational, for infinitely many $n$, $\tilde{c_{1}}n$, $\tilde{c_{2}}n$ and $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ are all integers. Furthermore, those $n$’s are a multiple of a given integer, depending on $r$.
Now we are going to construct a degree sequence $D^{\prime}$ for any degree sequence $D$. Let $S$ be the sum of the degrees in $D$. We might assume that $S$ is even, otherwise $D$ is surely not graphic. Further, we might assume that there is no degree which is at least $\frac{S}{2}$. If there is a degree which is larger than $\frac{S}{2}$, then $D$ cannot have a bipartite graph realization. Same is true if $D$ has more than one degree $\frac{S}{2}$ (except the trivial case $D=(1,1)$). Finally, if $D$ has exactly one degree $\frac{S}{2}$, then $D$ has a bipartite graph realization. We also might assume that $D$ has no 0 degrees, since 0 degree vertices can be freely moved from one vertex class to the other vertex class in a bipartite graph.
Let $n$ be the smallest integer for which $\tilde{c_{1}}n$, $\tilde{c_{2}}n$ and $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ are all integers, $\tilde{c_{1}}n$ and $\tilde{c_{2}n}$ are both multiple of $\frac{S}{2}$ and also
(11) $\frac{\tilde{c_{2}}n+\frac{S}{2}}{n+S}<c_{2}$
and
(12) $\frac{\tilde{c_{1}}n}{n+S}>c_{1}.$
Such an $n$ exists since $\tilde{c_{1}}>c_{1}$ and $\tilde{c_{2}}<c_{2}$. Further, this $n$ is a linear function of $S$ (and thus, a polynomial function of $|D|$). Indeed, the inequalities in equations 11 and 12 holds for
$n\geq\max\left\{\frac{S\left(\frac{1}{2}-c_{2}\right)}{c_{2}-\tilde{c_{2}}},\frac{Sc_{1}}{\tilde{c_{1}}-c_{1}}\right\},$
and we require that $n$ be a multiple of an integer depending on $r$ and $S$. Then the degree sequence $D^{\prime}$ contains $(\tilde{c_{1}}+\tilde{c_{2}})n$ instances of degree $\tilde{c_{2}}n+\frac{S}{2}$, for each $d_{i}\in D$, a degree $d_{i}+\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$, S|D|S-|D| instances of degree $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ and $n-(\tilde{c_{1}}+\tilde{c_{2}})n$ instances of degree $\tilde{c_{1}}n$. That is, $D^{\prime}$ has length $n+S$, and due to equations 11 and 12, $D^{\prime}$ is a problem instance in $\textsc{BDR}_{c_{1},c_{2}}$. Further, it can be constructed from $D$ in polynomial time.
We are going to prove that $D^{\prime}$ has a bipartite graph realization if and only if $D$ has a bipartite graph realization.
(Forward direction) That is, if $D$ has a bipartite graph realization, then $D^{\prime}$ also has a bipartite graph realization Let $G$ be a bipartite graph realization of $D$. Amalgamate $G$ with 0 degrees to obtain $\frac{S}{2}$ vertices in both vertex classes, let these vertex classes be $U$ and $V$. Add $\frac{n}{2}$ further vertices to both $U$ and $V$, identify among these added vertices $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ vertices in both vertex classes, and denote them by $U_{L}$ and $V_{L}$, and let the remaining $\frac{n}{2}-\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ vertices denoted by vertex classes $U_{S}$ and $V_{S}$. Create a complete bipartite graph between $U_{L}$ and $V_{L}$, similarly, create a complete bipartite graph between $U_{L}$ and $V$ and between $V_{L}$ and $U$. Then create a semi-regular bipartite graph between $U_{L}$ and $V_{S}$ such that in this semi-regular graph, the degrees in $U_{L}$ are $\frac{\tilde{c_{2}}-\tilde{c_{1}}}{2}n$, and the degrees in $V_{S}$ are $\tilde{c_{1}}n$. This is doable, since the sum of the degrees in the two vertex classes are the same:
$\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n\frac{\tilde{c_{2}}-\tilde{c_{1}}}{2}n=\frac{1-\tilde{c_{2}}-\tilde{c_{1}}}{2}n\tilde{c_{1}}n.$
Indeed, this is equivalent with $\tilde{c_{1}}=1-\tilde{c_{2}}-\sqrt{1-2\tilde{c_{2}}}$, according to Observation 4. Similarly, create a semi-regular graph between $V_{L}$ and $U_{S}$ with equivalent degrees. The so-obtained graph is a bipartite graph realization of $D^{\prime}$. Indeed, the degrees in $U_{L}$ (in $V_{L}$, respectively) are the sum of the degrees obtained from the two complete bipartite graphs towards $V_{L}$ and $V$ (to $U_{L}$ and $U$, respectively) and the degrees towards $U_{S}$ (respectively, $V_{S}$):
$\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}+\frac{\tilde{c_{2}}-\tilde{c_{1}}}{2}n=\tilde{c_{2}}n+\frac{S}{2}.$
Similarly, we can check that the degrees in $U$, $V$, $U_{S}$ and $V_{S}$ are the prescribed ones.
(Backward direction) That is, if $D^{\prime}$ has a bipartite graph realization, then $D$ also has a bipartite graph realization. Let $G^{\prime}=(U,V,E)$ be a bipartite graph realization of $D^{\prime}$. First, we claim that the degrees in form $d_{i}+\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ has the same sum in $U$ and in $V$. Indeed, the sum of the degrees in $U$ and $V$ must be the same, and thus, they must be the same mod $\frac{S}{2}$. However, each degree in $D^{\prime}$ is a multiple of $\frac{S}{2}$ except the degrees in form $d_{i}+\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$. Therefore, their sum also must equal in $U$ and $V$.
We further claim that both $U$ and $V$ contain $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ instances of degree $\tilde{c_{2}}n+\frac{S}{2}$. Assume contrary that $U$ has $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t$ instances of degree $\tilde{c_{2}}n+\frac{S}{2}$ and $V$ has $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t$ instances of degree $\tilde{c_{2}}n+\frac{S}{2}$, and w.l.o.g., $t>0$. First observe that the sum of the degrees in $D^{\prime}$ is
$(\tilde{c_{1}}+\tilde{c_{2}})n\left(\tilde{c_{2}}n+\frac{S}{2}\right)+(1-\tilde{c_{1}}-\tilde{c_{2}})n\tilde{c_{1}}n+S\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+S.$
Therefore, if $U$ contains $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t$ instances of degree $\tilde{c_{2}}n+\frac{S}{2}$, then the sum of those degrees in $V$ which are not $\tilde{c_{2}}n+\frac{S}{2}$ is
$\frac{(\tilde{c_{1}}+\tilde{c_{2}})n\left(\tilde{c_{2}}n+\frac{S}{2}\right)+(1-\tilde{c_{1}}-\tilde{c_{2}})n\tilde{c_{1}}n+S\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+S}{2}$-
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)\left(\tilde{c_{2}}n+\frac{S}{2}\right)=t\left(\tilde{c_{2}}n+\frac{S}{2}\right)+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}{n}+\frac{S}{2}\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}$
It follows that $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t<\tilde{c_{2}}n$, since otherwise the indicated sum could not be obtained as a sum of those degrees in $D^{\prime}$ which are not $\tilde{c_{2}}n+\frac{S}{2}$.
We claim that this split does not satisfy the Gale-Ryser inequality for the critical $k$ containing the degrees $\tilde{c_{2}}n+\frac{S}{2}$ and $d_{i}+\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ on the left-hand side. Assume that there are $x$ degrees in form $d_{i}+\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$. Then the left-hand side of the Gale-Ryser inequality is
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t\right)\left(c_{2}n+\frac{S}{2}\right)+x\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}$
Since $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t<\tilde{c_{2}}n$, it also holds that $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t+x<\tilde{c_{2}}n+\frac{S}{2}$ (for we assumed that $D$ has no 0 degrees, therefore $x\leq\frac{S}{2}$). Therefore, the right-hand side of the Gale-Ryser inequality upper bounded by
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t+x\right)+$
$t\left(\tilde{c_{2}}n+\frac{S}{2}\right)+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}n+\frac{S}{2}\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}=$
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n\right)^{2}-t^{2}+x\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)+$
$t\left(\tilde{c_{2}}n+\frac{S}{2}\right)+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}n+\frac{S}{2}\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}$
According to Observation 4, we know that
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n\right)^{2}+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}n=\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n\tilde{c_{2}}n,$
therefore the upper bound of the right-hand side of the Gale-Ryser inequality becomes
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t\right)\left(\tilde{c_{2}}n+\frac{S}{2}\right)-t^{2}+x\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)+\frac{S}{2}$
Subtracting this upper bound from the left-hand side we get
$t^{2}+xt,$
which is positive. Therefore in any realization $G=(U,V,E)$ of $D^{\prime}$, there are $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ instances of degree $\tilde{c_{2}}+\frac{S}{2}$ both in $U$ and in $V$. It follows that the left-hand side and the right-hand side are the same in the $k^{\mathrm{th}}$ Gale-Ryser inequality with $k=\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$. Further, it follows, if we remove the edges incident to vertices having degree $\tilde{c_{2}}+\frac{S}{2}$, the remaining graph will be a realization of $D$ amalgamated with several further vertices having degree 0. Therefore, $D$ has a bipartite graph realization.